Mr(Cu SO4 5HO)=1×Ar(Cu)+1×Ar(s)+9×Ar(o)+10×Ar(H)=?
1
Expert's answer
2017-11-27T11:22:07-0500
A_r (CuSO_4×5H_2 O)=1×A_r (Cu)+1×A_r (S)+9×A_r (O)+10×A_r (H)=(1*63.55+1*32.06+9*16.00+10*1.01) g⁄mol=249.71 g⁄mol.
Answer: A_r=(CuSO_4×5H_2 O)=249.71 g⁄mol.